If the delta G' value for an isomerization reaction, A --> A' is 8.0 kJ/mole at 37 degrees CBCH351 Rec. 304 Problem set 1-1 Due Jan. 18, 2018 Name: These problems worth 2 points (of the bonus points) J/mole at 37 C, a) what is the ratio of A and A at equilibrium? b) What is AG for the reaction if A is reacted soon after it is formed and its concen reduced to 1/50 x the normal equilibrium concentration? The concentration of A doesnt change and just consider the concentration effect on ΔG, (not linking the value of AG, of the second reaction onto this isomerization reaction) IS 2. The rate constant for the reaction depends on the Gibbs activation energy AG and the temperature (in degrees Kelvin) according to the equation: ke 2.08 x 1010 sKi *T*e(-AGキ/RT) a) If the rate constant for a reaction is 1.55 x 102 s-1 at 37 °C, what is the value of AG # for the reaction? b) What is the rate constant for the reverse reaction (same reaction as in problem 1)? 3. What are the pHs of solutions that have H concentrations of (a) 2.00 x 10 8 mol/L; (b) 4.0x 103 mol/L, (c) 2 mol/L?
Answer #1

1)

a)Given Data :

\DeltaG = 8000 J/mol

R = 8.314 J/mol K

T = 37 C = 310 K

\DeltaG = -RTlnK

lnK = - 8000/ 8.314 * 310 = -3.1039

K = 0.0448 = 0.045

K, equillibrium constant is nothing but ratio of concentration of products to concentration of reactants. So we have,

K = [A']/[A] = 0.045

b) [A'] becomes 1/50 times its value in the second case. K the becomes,

K = {[A']/50}/[A] = 0.045/50 = 0.0009

\DeltaG = -RTlnK

= - 8.314 * 310 * ln(0.0009)

= 18075.18 J/mol = 18.075 KJ/mol